The units of an item available for sale during the year were as follows:
Jan. 1 Inventory 27 units at $120
Feb. 17 Purchase 54 units at $138
July 21 Purchase 63 units at $156
Nov. 23 Purchase 36 units at $165
There are 50 units of the item in the physical inventory at December 31. The periodic
inventory system is used. Determine the inventory cost by (a) the first-in, first-out
method, (b) the last-in, first-out method, and (c) the average cost method.
1
Expert's answer
2014-11-13T10:17:47-0500
(a) According to the first-in, first-out method: Cost sold= 27*120 + 54*138 + 49*156 = $18,336
Ending inventory = 14*156 + 36*165 = $8,124
(b) According to the last-in, first-out method: Cost sold = 36*165 + 63*156 + 31*138 = $20,046
Ending inventory = 23*138 + 37*120 = $7,614
(c) According to the average cost method: Average cost = (27*120 + 54*138 + 63*156 + 36*165)/(27 + 54 + 63 + 36) = $147
Cost sold = 130*147 = $19,110
Ending inventory = 50*147 = $7,350